Suppose that U&I Inc. is a Canadian company that every year sells $10B of product, $6B to domestic and $4B to foreign consumers. Yearly, it buys $1B of imported parts, pays its workers $5.5B, pays rent of $0.5B, makes a $2B interest payment, and earns profits of $1B. U&I’s operations add $9B yearly to GDP whether measured using the value added, expenditure, or income approach. True, False, or Uncertain? Explain.


1
Expert's answer
2021-02-15T17:50:23-0500

Solution:

True.

Value-added approach:

Gross Value Added = Gross Value of Output – Value of Intermediate Consumption

GDP = Sum of all value-added to products during the production of a process

GDP = VOGS – IC

 = 10 – 1

= $9B


Expenditure approach:

GDP = C + G + I + NX

= 6 + 0 + 2 + (4 -1)

 = $9B

 

Income approach:

GDP = Total National Income + Sales Taxes + Depreciation + Net Foreign Factor Income

Total National Income – the sum of all wages, rent, interest, and profits.

 = 5.5 + 0.5 + 2 + 1

 = $9B